# Rethinking Circuit Completeness in Language Models: AND, OR, and ADDER Gates

### v1

@@ D.1.1 Greedy Search Example: ACDC
.$\mathcal{H}\leftarrow\mathcal{H}.{reverse\_topological\_sort()}$
@@
D.1.2 Linear Estimation Example: EAP
@@
$L(\tilde{x}|do(x_{i}))-L(\tilde{x})\approx(\tilde{x}_{i}-x_{i})^{T}\frac{\partial}{\partial\tilde{x}_{i]}}L(\tilde{x})$ (5)
Therefore, the approximation for Ns.+Dn. is $(\tilde{x}_{i}-x_{i})^{T}\frac{\partial}{\partial x_{i]}}L(x)+(\tilde{x}_{i}-x_{i})^{T}\frac{\partial}{\partial\tilde{x}_{i]}}L(\tilde{x})$.
D.1.3 Differentiable Mask Example: EdgePruning
@@
Appendix E Misalignment Score
@@
Equation 7 utilizes the properties that the effect does not change by removing one edge from OR gates, while it significantly changes from ADDER gates. Additionally, to avoid the bias caused by both effects of ADDER and OR edges being marginally small in large-scale circuits, we replace the “difference in one edge” with the “difference in difference between one edge and two edges,” and introduce a constant $m$ to ensure that the score $>0$ (with $m$ set to $1.5$ in practice).
Therefore, for any pair of $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$, we can compute the misalignment using these two scores. We report the misalignment scores resulting from the intersection of Ns and Dn circuits at varying scales. Specifically, we select an Ns circuit consisting of 100 edges recovered from the IOI task and examine how the misalignment score changes as the number of edges in the Dn circuit varies from 60 to 140. Figure 8 illustrates that when Dn is significantly smaller than Ns, the misalignment score for the AND gates is high, as many ADDER edges are misclassified as AND edges. Conversely, when Dn is substantially larger than Ns, the misalignment score for the OR gates increases, due to many ADDER edges being misclassified as OR edges. When the number of edges in the Dn circuit is approximately equal to that in the Ns circuit, both the misalignment score and its standard deviation reach an acceptable level. Therefore, throughout this paper, we assume that the optimal alignment occurs when Ns and Dn contain an equal number of edges.
@@
Appendix F Experiments of Faithfulness
@@
Appendix G Graph Study
@@
Figure 10: Circuit Graphs of AND, OR, and ADDER gates, respectively. We set the color of each component to be the same as that of the IOI circuit [Wang et al., ], allowing for easy reference to the function of each component.
To investigate the relationship between the three logical gates and the functions of the language model, we extracted the circuits of AND, OR, and ADDER gates discovered through ACDC on the IOI task. We then reviewed the IOI circuit [Wang et al., ] to determine the function of each component.
Interestingly, each receiver in the AND circuit is almost always influenced by edges from different functions, indicating that the AND operation can be understood as combining different functions to jointly impact the subsequent layers. For example, in Figure 10(a), the Induction Head requires edges from both the Duplication Token Head and the Previous Token Head to function, which supports the mechanism behind the Induction skill [Crosbie and Shutova, 2024, Ren et al., 2024, Edelman et al., 2024]. Similarly, the Name Mover Head requires support from both the S-Inhibition Head and the Induction Head, which explains the functional mechanism of the AND operation.
@@
Appendix H Output Contribution
@@
Appendix I Proportion of AND, OR, and ADDER Gates
@@
Figure 12 illustrates the proportion of the three types of logical gates across different tasks for each method. Notably, in the ACDC baseline, the number of edges corresponding to each gate type is nearly equal. This is because ACDC employs a greedy search strategy without ranking edges by their effect on the output. In contrast, both EAP and EdgePruning yield significantly fewer OR edges, reflecting the fact that OR edges contribute the least to the output—a finding we detail in Section 5.2 and Appendix H. Furthermore, the results from EdgePruning indicate that the number of AND edges is similarly low, comparable to OR edges. This arises from the fact that EdgePruning optimizes based on individual edge effects rather than gate-level effects. For instance, in a gate comprising two AND edges, each edge contributes only half of the total gate effect. As a result, during optimization, the mask values for such edges may be suppressed, increasing the likelihood of pruning.

### v2

@@ E.3.1 Greedy Search Example: ACDC
$\mathcal{H}\leftarrow\mathcal{H}.{reverse\_topological\_sort()}$
@@
E.3.2 Linear Estimation Example: EAP
@@
[…]
E.3.3 Differentiable Mask Example: EdgePruning
@@
Appendix F Misalignment Score
@@
Equation 7 utilizes the properties that the effect does not change by removing one edge from OR gates, while it significantly changes from ADDER gates. Additionally, to avoid the bias caused by both effects of ADDER and OR edges being marginally small in large-scale circuits, we replace the “difference in one edge” with the “difference in difference between one edge and two edges,” and introduce a constant $m$ to ensure that the score >0>0 (with $m$ set to $1.5$ in practice).
Therefore, for any pair of $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$, we can compute the misalignment using these two scores. We report the misalignment scores resulting from the intersection of Ns and Dn circuits at varying scales. Specifically, we select an Ns circuit consisting of 100 edges recovered from the IOI task and examine how the misalignment score changes as the number of edges in the Dn circuit varies from 60 to 140. Figure 8 illustrates that when Dn is significantly smaller than Ns, the misalignment score for the AND gates is high, as many ADDER edges are misclassified as AND edges. Conversely, when Dn is substantially larger than Ns, the misalignment score for the OR gates increases, due to many ADDER edges being misclassified as OR edges. When the number of edges in the Dn circuit is approximately equal to that in the Ns circuit, both the misalignment score and its standard deviation reach an acceptable level.
As shown in Figure 8, when the OR set significantly exceeds the AND set, a large number of misclassified ADDER gates appear only within the OR set, leading to a much higher misalignment score for OR compared to AND. The misalignment score is more of a "property," reflecting the optimal scale at which OR and AND sets should intersect. We determine the optimal ratio of AND to OR gates by minimizing the sum of the misalignment scores for AND and OR. The optimal results obtained with different methods and datasets are approximately as shown in Table 6:
Table 6: The optimal ratio of edge numbers between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$.
strategies IOI GT SA
ACDC 1.12: 1 1.05: 1 1.07: 1
EAP 1.09: 1 1.03:1 1.05: 1
Edge-Pruning 1.11: 1 1.07: 1 1.06: 1
The results show that the proportion of AND is always slightly higher than that of OR, but it can be approximated as 1:1. Therefore, throughout this paper, we assume that the optimal alignment occurs when Ns and Dn contain an equal number of edges.
@@
Appendix G Experiments of Completeness
In this Appendix, we present the detailed results of the completeness validation, together with a comparison against the Dn strategy. The specific results are reported in Tables 7 and  8. In brief, our method generally outperforms the approach based on Dn. Although the Dn-based method is theoretically capable of identifying all OR gates, it suffers from certain shortcomings in practice. For instance, as shown in Table 1, EAP fails to identify one of the edges associated with an AND gate under the Dn framework. Additionally, the greedy strategy may miss some OR gates located beneath branches of AND gates due to its dependence on the search order. The differentiable mask approach, on the other hand, suffers from the drawback that, in gates involving multiple edges, the mask values assigned to each edge tend to be lower, reducing its effectiveness.
Table 7: KL in Completeness Validation of IOI task, including Ns, Dn, and Ns+Dn.
#edges ACDC_NS ACDC_Dn ACDC_NsDn EAP_Ns EAP_Dn EAP_NsDn EdgeP_Ns EdgeP_Dn EdgeP_NsDn
100 0.62±0.1 0.64±0.1 0.64±0.1 0.51±0.1 0.71±0.1 0.74±0.1 0.56±0.1 0.62±0.1 0.64±0.1
200 0.64±0.1 0.69±0.1 0.71±0.1 0.56±0.1 0.73±0.0 0.73±0.1 0.58±0.1 0.65±0.1 0.69±0.1
500 0.71±0.1 0.84±0.1 0.87±0.1 0.65±0.1 0.81±0.1 0.83±0.1 0.67±0.1 0.88±0.1 0.94±0.1
1000 0.81±0.1 0.96±0.1 1.03±0.1 0.74±0.1 0.96±0.1 1.02±0.1 1.08±0.1 1.11±0.1 1.13±0.1
2000 1.05±0.1 1.19±0.2 1.22±0.1 0.89±0.1 1.11±0.2 1.24±0.2 1.31±0.2 1.40±0.1 1.41±0.1
5000 1.34±0.2 1.42±0.2 1.46±0.2 1.21±0.1 1.42±0.2 1.48±0.2 1.62±0.2 1.73±0.1 1.73±0.1
Table 8: Accuracy in Completeness Validation of IOI task, including Ns, Dn, and Ns+Dn.
#edges ACDC_NS ACDC_Dn ACDC_NsDn EAP_Ns EAP_Dn EAP_NsDn EdgeP_Ns EdgeP_Dn EdgeP_NsDn
100 0.73±0.1 0.68±0.0 0.63±0.1 0.74±0.1 0.53±0.0 0.51±0.0 0.76±0.1 0.67±0.1 0.66±0.0
200 0.65±0.0 0.62±0.0 0.60±0.0 0.68±0.0 0.49±0.0 0.46±0.0 0.59±0.0 0.55±0.0 0.56±0.0
500 0.53±0.0 0.51±0.0 0.51±0.0 0.61±0.0 0.41±0.0 0.38±0.0 0.55±0.0 0.43±0.0 0.43±0.0
1000 0.40±0.0 0.34±0.0 0.33±0.0 0.39±0.0 0.29±0.0 0.26±0.0 0.32±0.0 0.27±0.0 0.26±0.0
2000 0.34±0.0 0.26±0.0 0.24±0.0 0.35±0.0 0.18±0.0 0.14±0.0 0.24±0.0 0.18±0.0 0.16±0.0
5000 0.26±0.0 0.22±0.0 0.21±0.0 0.29±0.0 0.13±0.0 0.09±0.0 0.21±0.0 0.11±0.0 0.11±0.0
Appendix H Experiments of Faithfulness
@@
Appendix I Graph Study
@@
Figure 10: Circuit Graphs of AND, OR, and ADDER gates, respectively. We set the color of each component to be the same as that of the IOI circuit [16], allowing for easy reference to the function of each component.
To investigate the relationship between the three logical gates and the functions of the language model, we extracted the circuits of AND, OR, and ADDER gates discovered through ACDC on the IOI task. We then reviewed the IOI circuit [16] to determine the function of each component.
Interestingly, each receiver in the AND circuit is almost always influenced by edges from different functions, indicating that the AND operation can be understood as combining different functions to jointly impact the subsequent layers. For example, in Figure 10(a), the Induction Head requires edges from both the Duplication Token Head and the Previous Token Head to function, which supports the mechanism behind the Induction skill [27, 28, 29]. Similarly, the Name Mover Head requires support from both the S-Inhibition Head and the Induction Head, which explains the functional mechanism of the AND operation.
@@
Appendix J Output Contribution
@@
Appendix K Proportion of AND, OR, and ADDER Gates
@@
Figure 12 illustrates the proportion of the three types of logical gates across different tasks for each method. Notably, in the ACDC baseline, the number of edges corresponding to each gate type is nearly equal. This is because ACDC employs a greedy search strategy without ranking edges by their effect on the output. In contrast, both EAP and EdgePruning yield significantly fewer OR edges, reflecting the fact that OR edges contribute the least to the output—a finding we detail in Section 5.2 and Appendix J. Furthermore, the results from EdgePruning indicate that the number of AND edges is similarly low, comparable to OR edges. This arises from the fact that EdgePruning optimizes based on individual edge effects rather than gate-level effects. For instance, in a gate comprising two AND edges, each edge contributes only half of the total gate effect. As a result, during optimization, the mask values for such edges may be suppressed, increasing the likelihood of pruning.
Appendix L Potential Contributions of Completeness: an Example of Model Unlearning
We have designed a toy task analogous to knowledge unlearning to demonstrate its actual effectiveness. Specifically, given a harmful-response datasets (PKU-SafeRLHF) [ji2023beavertails], we first obtain its circuit $\mathcal{C}$, and then remove this circuit from the computation graph $\mathcal{G}$, observing whether the model can still respond to the dataset. The dataset provides harmful responses for certain questions, which are used to evaluate the efficacy of preventing the model from generating these harmful responses.
In the Table 9, we present the performance of this toy task on the accuracy of $\mathcal{G-C}$, which refers to the performance of the computational graph after the circuit has been removed (using the interchange ablation method), reflecting the completeness of the circuit. We test two circuit discovery methods: EAP and Edge-Pruning.
Table 9: A toy experiment on model unlearning task
Method accuracy accuracy of G-C
EAP 92.74 29.17
Edge-P 92.74 31.44
Ours 92.74 13.21
It is evident that, compared to existing circuit discovery methods, our method ensures that the computational graph loses almost all the necessary mechanisms for the task once the circuit is removed. Additionally, this toy task highlights the significance of studying circuit completeness, which holds great potential in unlearning harmful information: a complete circuit can help us pinpoint all neurons associated with harmful responses.